# Kartik Narayan, Vibashan VS, Vishal M. Patel

### v1

@@ 1 Introduction
Before exploring MLLMs for face understanding capabilities, it’s essential to address the fundamental question: “Why should MLLMs be proficient in face understanding?” MLLMs are increasingly deployed as central processors in various advanced applications, including virtual-reality headsets [38], embodied AI [29, 20], driving safety [31, 89], authentication [18], human-computer interaction [100], and sports analysis [106]. In these applications, accurate face understanding is crucial, as face images appear frequently and require accurate face understanding for appropriate responses. However, the face understanding capabilities of existing MLLMs are limited; they often fail to answer basic questions such as “What is the expression of the person in this image?” or “Which of the following regions is not present in the face image?” These shortcomings indicate a significant scope for improvement. Recent works, such as EMO-LLaMA [109], develop an instruction-tuning set for supervised fine-tuning to enhance MLLMs’ capabilities in understanding facial expressions. Face-MLLM [91] proposed a three-stage training pipeline to equip MLLMs’ with face perception capabilities. Nonetheless, the research community currently lacks a standardized benchmark to quantify and compare MLLMs’ performance in face understanding. We believe that a comprehensive benchmark incorporating various aspects of face understanding is a crucial and essential foundational step for monitoring progress and advancing MLLMs’ performance in this domain.
@@
Recent works [75, 93] emphasize that, although existing open-source MLLMs’ have achieved advanced capabilities, they still lack the sophistication to perform complex tasks. To address this, these approaches enable MLLMs with tool use, extending their capabilities by leveraging external tools to fulfill complex human instructions that require task-specific processing. Motivated by this, we introduce FaceXAPI, a new dataset that forms part of FaceXBench. It is designed to evaluate MLLMs’ ability to select the appropriate API and function calls for handling complex tasks in face understanding. FaceXBench comprises $5,000$ carefully and manually filtered VQA-type questions derived from $25$ public datasets, covering $14$ tasks. It includes $10,441$ unique face images representing a diverse range of age groups, genders, races, varied resolutions, head poses, expressions and attributes, reflecting the diversity of faces encountered in real-world scenarios.
@@
  • Extensive Evaluation: We evaluate $26$ open-source MLLMs and two proprietary models, GPT-4o and GeminiPro1.5, which achieve accuracies of $50.24$% and $54.40$%, respectively, underscoring the significant challenge posed by FaceXBench and the substantial room for improvement.
@@ 3 The FaceXBench Benchmark
Table 1: Key statistics of question in FaceXBench.
@@
3.1 Overview of FaceXBench
We introduce FaceXBench, a novel and comprehensive benchmark encompassing multiple aspects of face understanding. Additionally, we propose FaceXAPI, a manually curated dataset within FaceXBench that aims to evaluate tool-use capabilities. FaceXBench is the first benchmark specifically designed to evaluate MLLMs’ performance on face-related tasks. It assess MLLMs’ across six broad categories: bias and fairness, face authentication, face recognition, face analysis, face localization, and facial tool use. Fig. 2 presents a detailed taxonomy of these categories, outlining their respective tasks and corresponding number of questions. The designed questions test a MLLMs’ capabilities in areas such as visual grounding, fine-grained feature extraction, anomaly detection, emotion recognition, fairness, contextual understanding, spatial understanding and agentic reasoning.
Benchmark Statistics. FaceXBench consists of $5,000$ questions, divided into $6$ categories covering $14$ tasks, and is derived from $25$ public datasets along with one proposed dataset (i.e. FaceXAPI). There are $2,750$ questions with multiple images, $2,150$ single-image questions, and $100$ text-only questions. The questions are generated using $757$ unique question templates and contain $10,441$ unique images. The correct answer options are approximately equally distributed between A, B, C and D to avoid bias. They key statistics of the benchmark are summarized in Tab. 2.
@@ 3.2 Data Collection
Our data collection pipeline consists of three steps. Step 1. In the first step, we iterate through the identified tasks in each category and select the datasets corresponding to each task. We collect the test sets of existing standard datasets for each task to avoid data leakage. We also created a new dataset, FaceXAPI, specifically for the tool-retrieval task. In total, our collection includes $25$ public datasets along with this newly proposed dataset. Step 2. In this step, we manually created question templates for each task involving single and multiple images. While creating these templates, we carefully framed questions to encourage the model to reason, compare, and think critically to find the correct answer. We included a mix of easy and hard questions to maintain a balanced distribution. For example, a hard question is, “<image1><image2><image3> How many images show a person in the age range of $30$ to $39$?,” while an easy question is, “<image1> What is the age range of the person shown in the image?”. We prompt GPT-4o with the manually curated question templates as in-context examples and generate additional templates that address similar questions but with varied phrasing, perspective, and/or complexity. For example, “<image1><image2> Which among these two appears to be older?”. We manually filter the question templates for each task to ensure diversity and task relevance. Additional details about the question templates and their generation can be found in Appendix C.1. Step 3. The final step involves generating answer options. Each question includes four options, with one correct answer. We ensure that distractor options are close to the correct answer but distinct enough to require critical thinking and reasoning when selecting the correct response. For example, for the question “<image1> What is the age range of the person shown in the image?”, the options designed are (A) $20$ to $29$, (B) $10$ to $19$ (C) $30$ to $39$ and (D) $40$ to $49$, with the correct answer being (C) $30$ to $29$. The thoughtfully chosen distractor options close to the actual answer, encourages the model to reason, compare and carefully select the correct choice. This approach raises the difficulty for the model, making the questions more challengeing. To validate this, we conducted a mini experiment using llava onevision qwen2 7b ov for age estimation on FairFace dataset questions that include single image. The model achieved $88$% accuracy with randomly chosen options, compared to $53$% when strategically designed options were used. This result highlights the importance of identifying effective distractor options. Further details on generating distractor options for each task are provided in Appendix C.2. Overall, we observed that the model’s performance significantly varies depending on the complexity of the questions and distractor options.
@@ 3.3 FaceXAPI
We believe that face understanding is an application domain that can be better addressed by equipping MLLMs with tools rather than relying solely on supervised fine-tuning. We provide a detailed discussion, supported by experimental validation, in the Discussion and Future Directions section (Section 6). To test MLLMs’ ability to select the correct sequence of APIs and function calls for successful task completion, we created FaceXAPI, a dataset of $100$ text-only questions referencing 13 APIs and 32 function calls. The questions were designed to ensure diversity in scenarios, reflecting a broad range of real-world applications and requiring a sequence of $3$ to $5$ API calls to solve. The dataset includes a total of 88 unique function call sequences. The questions were generated using GPT-4o, followed by manual filtering. A sample from the dataset is presented in Fig. 3, with the complete prompt provided in Appendix C.3.
@@ 4 Experiments
In this section, we detail the experiments conducted to benchmark and analyze various MLLMs on face understanding. We evaluate $2$ proprietary and $26$ open-source models, as listed in Section 4.1. For a fair comparison, all models are evaluated in a zero-shot setting using the same base prompt. We analyze selected MLLMs under in-context and chain-of-thought settings. The evaluation settings are described in Section 4.2 and the results are presented in Section 5. All experiments are performed using 8 NVIDIA A6000 GPU cards.
@@ 4.1 Models
The $2$ proprietary models used are GPT-4o [30] and GeminiPro 1.5 [96]. We divide the $26$ open-source models into three major categories based on parameter size: (a) Open Source MLLMs (<4B parameters): PaliGemma [9], LLaVA-OneVision-0.5b-OV [47], and VILA 1.5-3b [54]; (b) Open Source MLLMs (4B-13B parameters): Chameleon-7b [95], Eagle-X4-8B-Plus [85], Idefics2-8b [41], Idefics-9b-Instruct [40], LLaVA-v1.5-7b [57], Monkey-Chat [53], MiniCPM-Llama3-v2.5 [114], LLaVA-OneVision-7b-SI [47], LLaVA-NeXT-Interleave-7b [48], Mantis-SIGLIP-8b [32], Phi-3.5-Vision [1], LLaVA-OneVision-7b-OV, Qwen2-VL-7b-Instruct [103], and InternVL2-8b [12]; (c) Open Source MLLMs (>13B parameters): CogVLM2-19b [26], Idefics-80b-Instruct [40], LLaVA-v1.5-13b [57], VILA 1.5-13b [54], InternVL-Chat-v1.5 [12], VILA 1.5-40b [54], LLaVA-OneVision-72b-OV [47], Qwen2-VL-72b-Instruct [103], and InternVL2-76b [12]. In Appendix D.1, we provide detailed information regarding the architecture and the parameter size for all open-source MLLMs’ evaluated in this paper, along with additional results under different settings.
@@ 4.2 Evaluation Settings
We evaluate the models under three settings: (a) zero-shot, (b) in-context task description, and (c) chain-of-thought prompting. In the zero-shot setting, we pass the input with only the base prompt. In the in-context task description setting, we prepend the prompt with a brief description of the specific task required by the question. In the chain-of-thought setting, we prompt the model to reason step-by-step before selecting the correct option. The task-specific prepended text and additional details of the evaluation settings are provided in Appendix C.6.
The proposed benchmark consists of multiple-choice questions (MCQs), which standardize the evaluation across different models. We empirically construct a diverse set of regular expressions and employ a three-step evaluation strategy to extract the correct option in cases where intermediate reasoning or calculations are present in the model’s output. Our evaluation strategy is as follows: First, we use a regular expression to match the correct option (A, B, C, or D) at the beginning of the model’s prediction. If this attempt fails, we then search for the correct option within the entire prediction text. As a final fallback, we compare parts of the predicted output with the option values to find a match. If none of these steps succeed, we categorize the prediction as incorrect. The complete function is provided in Appendix C.4. Alternatively, we could implement a random-choice or frequent-choice strategy instead of labeling responses as incorrect. In the random-choice approach, a random option is selected as the prediction, while in the frequent-choice approach, we select the most common correct option in the dataset. However, these strategies can lead to misleading results, especially when the model fails to respond due to content moderation measures or poor instruction-following.
@@ 5 Results
Frequent Choice 32.22 30.73 29.50 40.14 33.25 29.73 40.00
@@
The performance of various models on FaceXBench is shown in Table 2. The results emphasize the challenging nature of the benchmark, with no model achieving more than $60$% accuracy. We observe that Qwen2-VL-72b-Instruct [103] achieves the best overall performance of 57.86%. InternVL2-76b [12], GeminiPro 1.5 [96], Qwen2-VL-72b-Instruct [103], LLaVA-OneVision-72b-OV [47], Qwen2-VL-72b-Instruct [103] and GeminiPro 1.5 [96] achieve the highest performances in the categories of Bias & Fairness, Face Recognition, Face Authentication, Face Analysis, Face Localization and Face Tools use with accuracies of $69.53$%, $70.00$%, $41.14$%, $63.25$%, $55.45$% and $57.00$%, respectively.
@@
Performance across various tasks. To gauge the difficulty of various tasks, we plot the performance of the top-5 models ($4$B-$13$B parameters) across each task, as shown in Figure 4(a). A detailed table showing the performance of all models across all tasks is provided in Appendix D.2 (Table D.2). We observe that, on average, models struggle with tasks such as crowd counting, deepfake detection, head pose estimation, and low-resolution face recognition, highlighting areas where MLLMs’ need improvement. In contrast, gender prediction is one of the easier tasks, with an average performance of approximately $80$%. Additionally, we plot the average performance of these top-5 models on multiple-image and single-image question. Figure 4(c) shows that models generally perform poorly on questions with multiple images as inputs compared to single-image questions. This is primarily because the models are required to process more visual information and compare and contrast facial features across multiple images, making the questions more challenging. Furthermore, models such as Qwen2-VL [103] and LLaVA-OneVision [47], which use dynamic resolution to handle arbitrary image resolutions, perform better than models on segmentation task. This mapping of images to a dynamic number of visual tokens more closely resembles human-like visual processing and leads to improved performance in face-understanding tasks.
@@ 7 Conclusion
1Table of Contents in Appendix
Appendix A Limitations

### v2

@@ 1 Introduction
Before exploring MLLMs for face understanding capabilities, it’s essential to address the fundamental question: “Why should MLLMs be proficient in face understanding?” MLLMs are increasingly deployed as central processors in various advanced applications, including virtual-reality headsets [39], embodied AI [30, 21], driving safety [32, 92], authentication [19], human-computer interaction [103], and sports analysis [109]. In these applications, accurate face understanding is crucial, as face images appear frequently and require accurate face understanding for appropriate responses. However, the face understanding capabilities of existing MLLMs are limited; they often fail to answer basic questions such as “What is the expression of the person in this image?” or “What is the age of the person in the face image?” These shortcomings indicate a significant scope for improvement. Recent works, such as EMO-LLaMA [112], develop an instruction-tuning set for supervised fine-tuning to enhance MLLMs’ capabilities in understanding facial expressions. Face-MLLM [94] proposed a three-stage training pipeline to equip MLLMs’ with face perception capabilities. Nonetheless, the research community currently lacks a standardized benchmark to quantify and compare MLLMs’ performance in face understanding. We believe that a comprehensive benchmark incorporating various aspects of face understanding is a crucial and essential foundational step for monitoring progress and advancing MLLMs’ performance in this domain. A more detailed discussion, addressing additional questions that frame our motivation, is provided in Appendix A.
@@
Recent works [78, 96] emphasize that, although existing open-source MLLMs’ have achieved advanced capabilities, they still lack the sophistication to perform complex tasks. To address this, these approaches enable MLLMs with tool use, extending their capabilities by leveraging external tools to fulfill complex human instructions that require task-specific processing. Motivated by this, we introduce FaceXAPI, a new dataset that forms part of FaceXBench. It is designed to evaluate MLLMs’ ability to select the appropriate API and function calls for handling complex tasks in face understanding. FaceXBench comprises $5,000$ carefully and manually filtered VQA-type questions derived from $25$ public datasets, covering $14$ tasks. Following established benchmarks such as MMBench [62] and SEEDBench [47], we curate multiple-choice questions derived from off-the-shelf datasets to maintain consistency and standardization, ensuring ease of use. In contrast, open-ended facial VQA relies on external models for answer matching, making the evaluation non-deterministic. Additionally, FaceXBench includes $10,441$ unique face images representing a diverse range of age groups, genders, races, varied resolutions, head poses and expressions, reflecting the diversity of faces encountered in real-world scenarios.
@@
  • Extensive Evaluation: We evaluate $26$ open-source MLLMs and two proprietary models, GPT-4o and GeminiPro1.5, which achieve accuracies of $50.24$% and $54.40$%, respectively, highlighting the significant challenge posed by FaceXBench and the substantial room for improvement.
@@ 3 The FaceXBench Benchmark
3.1 Overview of FaceXBench
We introduce FaceXBench, a novel and comprehensive benchmark encompassing multiple aspects of face understanding. Additionally, we propose FaceXAPI, a manually curated dataset within FaceXBench that aims to evaluate tool-use capabilities. FaceXBench is the first benchmark specifically designed to evaluate MLLMs’ performance on face-related tasks. It assess MLLMs’ across six broad categories: bias and fairness, face authentication, face recognition, face analysis, face localization, and facial tool use. Figure 2 presents a detailed taxonomy of these categories, outlining their respective tasks and corresponding number of questions. The designed questions test a MLLMs’ capabilities in areas such as visual grounding, fine-grained feature extraction, anomaly detection, emotion recognition, fairness, contextual understanding, spatial understanding and agentic reasoning.
Benchmark Statistics. FaceXBench consists of $5,000$ questions, divided into $6$ categories covering $14$ tasks, and is derived from $25$ public datasets along with one proposed dataset (i.e. FaceXAPI). There are $2,750$ questions with multiple images, $2,150$ single-image questions, and $100$ text-only questions. The questions are generated using $757$ unique question templates and contain $10,441$ unique images. The correct answer options are approximately equally distributed between A, B, C and D to avoid bias. They key statistics of the benchmark are summarized in Table 1.
Table 1: Key statistics of questions in FaceXBench.
@@ 3.2 Data Collection
Our data collection pipeline consists of three steps. Step 1. In the first step, we iterate through the identified tasks in each category and select the datasets corresponding to each task. We collect the test sets of existing standard datasets for each task to avoid data leakage. We also created a new dataset, FaceXAPI, specifically for the tool-retrieval task. In total, our collection includes $25$ public datasets along with this newly proposed dataset. Step 2. In this step, we manually created question templates for each task involving single and multiple images. While creating these templates, we carefully framed questions to encourage the model to reason, compare, and think critically to find the correct answer. We included a mix of easy and hard questions to maintain a balanced distribution. For example, a hard question is, “<image1><image2><image3> How many images show a person in the age range of $30$ to $39$?,” while an easy question is, “<image1> What is the age range of the person shown in the image?”. We prompt GPT-4o with the manually curated question templates as in-context examples and generate additional templates that address similar questions but with varied phrasing, perspective, and/or complexity. For example, “<image1><image2> Which among these two appears to be older?”. We manually filter the question templates for each task to ensure diversity and task relevance. Additional details about the question templates and their generation can be found in Appendix D.1. Step 3. The final step involves generating answer options. Each question includes four options, with one correct answer. We ensure that distractor options are close to the correct answer but distinct enough to require critical thinking and reasoning when selecting the correct response. For example, for the question “<image1> What is the age range of the person shown in the image?”, the options designed are (A) $20$ to $29$, (B) $10$ to $19$ (C) $30$ to $39$ and (D) $40$ to $49$, with the correct answer being (C) $20$ to $29$. The thoughtfully chosen distractor options close to the actual answer, encourages the model to reason, compare and carefully select the correct choice. This approach raises the difficulty for the model, making the questions more challenging. To validate this, we conduct a mini experiment using LLaVA-OV for age estimation on FairFace dataset questions. The model achieved $88$% accuracy with randomly chosen options, compared to $53$% when strategically designed options were used. This result highlights the importance of identifying effective distractor options. Further details on generating distractor options for each task are provided in Appendix D.2. Overall, we observed that the model’s performance significantly varies depending on the complexity of the questions and distractor options.
@@ 3.3 FaceXAPI
We believe that face understanding is an application domain that can be better addressed by equipping MLLMs with tools rather than relying solely on supervised fine-tuning. We provide a detailed discussion, supported by experimental validation, in the Discussion and Future Directions section (Section 6). To test MLLMs’ ability to select the correct sequence of APIs and function calls for successful task completion, we created FaceXAPI, a dataset of $100$ text-only questions referencing 13 APIs and 32 function calls. The questions were designed to ensure diversity in scenarios, reflecting a broad range of real-world applications and requiring a sequence of $3$ to $5$ API calls to solve. The dataset includes a total of 88 unique function call sequences. The questions were generated using GPT-4o, followed by manual filtering. A sample from the dataset is presented in Figure 3, with the complete prompt provided in Appendix D.3.
@@ 4 Experiments
In this section, we detail the experiments conducted to benchmark and analyze various MLLMs on face understanding. We evaluate $2$ proprietary and $26$ open-source models, listed in Section 4.1. For a fair comparison, all models are evaluated in a zero-shot setting using the same base prompt. We analyze selected MLLMs under in-context and chain-of-thought settings. The evaluation settings and results are described in Sections 4.2 and 5, respectively.
@@ 4.1 Models
The $2$ proprietary models used are GPT-4o [31] and GeminiPro 1.5 [99]. We divide the $26$ open-source models into three major categories based on parameter size: (a) Open Source MLLMs (<4B parameters): PaliGemma [9], LLaVA-OneVision-0.5b-OV [49], and VILA 1.5-3b [56]; (b) Open Source MLLMs (4B-13B parameters): Chameleon-7b [98], Eagle-X4-8B-Plus [88], Idefics2-8b [42], Idefics-9b-Instruct [41], LLaVA-v1.5-7b [59], Monkey-Chat [55], MiniCPM-Llama3-v2.5 [117], LLaVA-OneVision-7b-SI [49], LLaVA-NeXT-Interleave-7b [50], Mantis-SIGLIP-8b [33], Phi-3.5-Vision [1], LLaVA-OneVision-7b-OV, Qwen2-VL-7b-Instruct [106], and InternVL2-8b [12]; (c) Open Source MLLMs (>13B parameters): CogVLM2-19b [27], Idefics-80b-Instruct [41], LLaVA-v1.5-13b [59], VILA 1.5-13b [56], InternVL-Chat-v1.5 [12], VILA 1.5-40b [56], LLaVA-OneVision-72b-OV [49], Qwen2-VL-72b-Instruct [106], and InternVL2-76b [12]. In the Appendix E.1, we provide detailed information regarding the architecture and the parameter size for all open-source MLLMs’ evaluated, along with additional results under different settings.
@@ 4.2 Evaluation Settings
We evaluate the models under three settings: (a) zero-shot, (b) in-context task description, and (c) chain-of-thought prompting. In the zero-shot setting, we pass the input with only the base prompt. In the in-context task description setting, we prepend the prompt with a brief description of the specific task required by the question. In the chain-of-thought setting, we prompt the model to reason step-by-step before selecting the correct option. The task-specific prepended text and additional details of the evaluation settings are provided in the Appendix D.6.
The proposed benchmark consists of multiple-choice questions (MCQs), which standardize the evaluation across different models. We empirically construct a diverse set of regular expressions and employ a three-step evaluation strategy to extract the correct option in cases where intermediate reasoning or calculations are present in the model’s output. Our evaluation strategy is as follows: First, we use a regular expression to match the correct option (A, B, C, or D) at the beginning of the model’s prediction. If this attempt fails, we then search for the correct option within the entire prediction text. As a final fallback, we compare parts of the predicted output with the option values to find a match. If none of these steps succeed, we categorize the prediction as incorrect. The complete function is provided in Appendix D.4. Alternatively, we could implement a random-choice or frequent-choice strategy instead of labeling responses as incorrect. In the random-choice approach, a random option is selected as the prediction, while in the frequent-choice approach, we select the most common correct option in the dataset. However, these strategies can lead to misleading results, especially when the model fails to respond due to content moderation measures or poor instruction-following.
@@ 5 Results
Frequent Choice 26.68 25.60 26.88 35.14 28.38 26.55 40.00
Human 70.28 72.33 65.50 66.00 76.12 67.27 94.00
Vision SOTA models 84.50 84.33 81.87 89.57 91.37 80.90 57.00
@@
The performance of various models on FaceXBench is shown in Table 2. The results emphasize the challenging nature of the benchmark, with no model achieving more than $60$% accuracy. We observe that Qwen2-VL-72b-Instruct [106] achieves the best overall performance of 57.86%. InternVL2-76b [12], GeminiPro 1.5 [99], Qwen2-VL-72b-Instruct [106], LLaVA-OneVision-72b-OV [49], Qwen2-VL-72b-Instruct [106] and GeminiPro 1.5 [99] achieve the highest performances in the categories of Bias & Fairness, Face Recognition, Face Authentication, Face Analysis, Face Localization and Face Tools Use with accuracies of $69.53$%, $70.00$%, $41.14$%, $63.25$%, $55.45$% and $57.00$%, respectively. Human performance at $70.28$% and SOTA vision models at $84.50$%, serve as the upper bounds for the benchmark. The significant gap between the current state-of-the-art MLLM and this upper bound highlights the substantial room for improvement.
@@
Performance across various tasks. To gauge the difficulty of various tasks, we plot the performance of the top-5 models ($4$B-$13$B parameters) across each task, as shown in Figure 4(a). A detailed table showing the performance of all models across all tasks is provided in Appendix E.2. We observe that, on average, models struggle with tasks such as crowd counting, deepfake detection, head pose estimation, and low-resolution face recognition, highlighting areas where MLLMs’ need improvement. In contrast, gender prediction is one of the easier tasks, with an average performance of approximately $80$%. Additionally, we plot the average performance of these top-5 models on multiple-image and single-image question. Figure 4(c) shows that models generally perform poorly on questions with multiple images as inputs compared to single-image questions. This is primarily because the models are required to process more visual information and compare and contrast facial features across multiple images, making the questions more challenging. Furthermore, models such as Qwen2-VL [106] and LLaVA-OneVision [49], which use dynamic resolution to handle arbitrary image resolutions, perform better than models on segmentation task. This mapping of images to a dynamic number of visual tokens more closely resembles human-like visual processing and leads to improved performance in face-understanding tasks.
@@ 7 Conclusion
1Table of Contents
Appendix A Motivation
In the main paper, we already discussed “Why should MLLMs be proficient in face understanding?”. To reiterate, MLLMs are increasingly deployed as central processors in various advanced applications, including virtual-reality headsets [39], embodied AI [30, 21], driving safety [32, 92], authentication [19], human-computer interaction [103], and sports analysis [109]. In these applications, accurate face understanding is crucial, as face images appear frequently and require accurate face understanding for appropriate responses.
A common question that arises is “Why do we need MLLMs to accomplish face analysis task, which are mostly perception-based, when MLLMs are not explicitly trained for these tasks?”. To clarify, our objective is not to solve face analysis tasks using MLLMs, nor do we aim to outperform specialized face analysis models on established benchmarks. Instead, our focus is on evaluating MLLMs’ ability to understand faces, which is a fundamental aspect of general visual and multimodal reasoning. Strong face understanding is essential for a variety of real-world applications, as discussed in the previous question. Given that MLLMs are expected to serve as general-purpose vision-language models, assessing their performance in this domain provides valuable insights into their limitations and potential areas for improvement. FaceXBench serves as a standardized benchmark to systematically evaluate and monitor the progress of MLLMs in face understanding. By identifying gaps and shortcomings in their capabilities, our work can help guide future advancements in MLLMs, ensuring that they develop stronger and more reliable face understanding skills.
Another important question that comes to mind is “Face analysis tasks are primarily fine-grained and vision-centric, requiring little to no language reasoning. Why is it necessary to evaluate MLLMs on face understanding rather than using specialized vision models for face analysis and then passing the results to LLMs?”. We acknowledge that using an external tool or agent leveraging specialized vision models and then passing the results to LLMs would likely yield better performance in face analysis tasks. To validate this, we have included tool-use results in Table 3(b) of the main paper. However, the primary goal of FaceXBench is not to optimize performance through external assistance but to directly assess the inherent face-understanding capabilities of existing MLLMs. As discussed in the introduction (lines 46–72) of main paper, we argue that MLLMs should possess strong face-understanding skills intrinsically, as this ability is crucial for a wide range of real-world multimodal applications. Many general tasks such as identity-aware reasoning, social interaction analysis, and facial emotion recognition require a model to process and reason about facial features without relying on a separate vision module. Evaluating MLLMs independently allows us to benchmark their progress, identify limitations, and drive improvements in their multimodal reasoning capabilities, ultimately making them more robust for downstream applications where face understanding is essential.
We feel that it also important to answer the question “What is the rationale behind our selected tasks? For example, why do we not include face detection or prompt MLLMs to generate face parsing maps?”. We select tasks that are most relevant to the capabilities and real-world applications of MLLMs. While face detection and direct face segmentation are important computer vision tasks, we exclude them from our benchmark because current MLLMs generally fail to handle them effectively and falls outside the intended use of MLLMs. Using a large MLLM to carry out face detection or parsing is thus impractical , instead, we include tasks such as crowd counting and evaluating segmented attributes because these tests still probe an MLLM’s ability to localize and reason about faces, which is essential for many downstream applications that MLLMs are designed to solve. Again, we would like to emphasize that our goal is not to solve face analysis tasks outright; rather, we select and adapt tasks that suit MLLMs and align with their purpose and intended use.
Appendix B Limitations